OpenStudy (anonymous):

What is the product of the complex numbers (-3i+4) and (3i+4) ?

OpenStudy (alexandervonhumboldt2):

(-3i+4)*(3i+4)=-9*(-1)-12i+12i+16=9+16=25

OpenStudy (jango_in_dtown):

(a+bi)(c+di)=(ac-bd)+i(ad+bc)

OpenStudy (alexandervonhumboldt2):

i^2=-1 remember

OpenStudy (anonymous):

thanks